HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

MICROWAVE SPECTRA OF THE CD_{3}OD and CH_{3}OD METHANOL ISOTOPOLOGS UP TO 1.1 THz AND THEIR SEARCH TOWARD IRAS

16293-2422

V. ILYUSHIN, \underline{E. ALEKSEEV} , R. POROHOVOI, O. DOROVSKAYA, Institute

of Radio Astronomy of NASU, Mystetstv 4, 61002 Kharkiv, Ukraine; \underline{H. S. P. MÜLLER} , F. LEWEN, S. SCHLEMMER, I. Physikalisches Institut,

Universität zu Köln, 50937 Köln, Germany; C. MAUL, S. BAUERECKER, Institut für Physikalische und Theoretische Chemie, Technische Universität Braunschweig, Gaußstr. 17, 38106 Braunschweig, Germany; J. K. JØRGENSEN, Niels Bohr Institute, University of Copenhagen, Øster Voldgade 5-7, 1350 Copenhagen K, Denmark; R. LEES, Department of Physics, University of New Brunswick, Saint

John, NB E2L 4L5, Canada

Different isotopic species of methanol have been detected in the warm parts of star-forming regions. We have started a program to investigate their rotational spectra to facilitate searches for these in space and here we present the results of our study of torsion-rotation spectra of the two deuterated isotopologs of methanol: CD_{3}OD and CH_{3}OD. The new microwave measurements were carried out using spectrometers in Kharkiv and Köln (in total from 34 GHz up to 1.1 THz). The analysis is done using the rho axis method and the RAM36 program code. For both isotopologs the v_{t} = 2 torsional state is significantly affected by intervibrational interactions with non-torsional vibrational modes which propagate down through intertorsional interactions. Thus we decided to concentrate our efforts on analysis of the v_{t} = 0, 1 states at the moment. For both isotopologs we were able to get a fit within experimental error for v_{t} = 0, 1 states and for both isotopologs we derived a line list for radio-astronomical observations. A search for CD_{3}OD in data from the Protostellar Interferometric Line Survey (PILS) of IRAS 16293-2422 obtained with ALMA was performed. While we found several emission features that can be attributed largely to CD_{3}OD, their number is as yet insufficient to establish a clear detection. For CH_{3}OD the analysis of PILS observational data is in progress.

The work in Cologne was supported by the Deutsche Forschungsgemeinschaft (DFG) via the collaborative research center SFB 956, project area B3, and through the Gerätezentrum “Cologne Center for Terahertz Spectroscopy”. The work in Kharkiv was supported by the Volkswagen Foundation. The assistance of the Science and Technology Center in the Ukraine is acknowledged (STCU partner project P756).

p-number: p125  Submitted on Wed, 14 Jun 2023 11:21:25 +0200